OpenStudy (clamin):

Which of the binomials below is a factor of this trinomial? x^2-13x+36 a. x+9 b. x+6 c. x-9 d. x-6

zepdrix (zepdrix):

So let's check out some factors of 36: 12*3 6*6 9*4 18*2 Is there any way we can get 13 by `adding` a pair of these? 12+3 = 15 Hmm that pair doesn't work.

OpenStudy (clamin):

im thinking x-6??

RhondaSommer (rhondasommer):

BRAVO!

zepdrix (zepdrix):

That doesn't really answer my question :\ So I'm not sure how you came to that conclusion.

OpenStudy (clamin):

i was thinking 6*6=36 and its negative

RhondaSommer (rhondasommer):

she is correct :)

RhondaSommer (rhondasommer):

oh wait no she isnt

jhonyy9 (jhonyy9):

x^2 -13x +36 =0 x*x-9x-4x+(-9)(-4)=0 hope these will help you

RhondaSommer (rhondasommer):

-6 +6 is not 13 :(

RhondaSommer (rhondasommer):

I think it might be c